### v4

@@ 3.2 Preconditioners for restricted KRR
where $\bm{\widehat{G}}\in{}^{k\times k}$ approximates the Gram matrix $\bm{G}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})$.
FALKON, the current state-of-the-art
@@
These entries are clearly linked together in expectation:
@@
The main limitation of FALKON-type preconditioners is the large size of the Monte Carlo approximation error. To account for the Monte Carlo error, the available analyses [44, 45] assume a large number of centers $k=\Omega(\sqrt{N})$ and a large regularization $\mu=\Omega(\sqrt{N})$. When these assumptions are satisfied, FALKON-type preconditioners solve the restricted KRR equations in $\mathcal{O}((N+k^{2})k)$ operations. However, the assumptions on $k$ and $\mu$ may not be satisfied in practice. For example, $\mu$ is often small when chosen by cross-validation, and [44] contains examples of $\mu$ values up to five orders of magnitude smaller than $\mu=\sqrt{N}$.
@@
1:Embedding dimension $d$, input dimension $N$, sparsity level $\zeta$.
2:Embedding matrix $\bm{\Phi}\in{}^{d\times N}$ in sparse (e.g., CSR) format.
3:$\mathtt{rows}\leftarrow\bm{0}_{\zeta N},\mathtt{cols}\leftarrow\bm{0}_{\zeta N},\mathtt{vals}\leftarrow\bm{0}_{\zeta N}$
@@
9:$\bm{\Phi}\leftarrow\textsc{Sparse}(\mathtt{rows},\mathtt{cols},\mathtt{vals},d,N)$triangleright $d\times N$ sparse matrix with specified entries
Block Lanczos preconditioner
A different preconditioner suggested by [25] approximates the Gram matrix $\bm{G}$ using the randomized block Lanczos algorithm [35]. Block Lanczos produces a high-quality rank-$r$ approximation in $\mathcal{O}(Nkr\log k)$ arithmetic operations. The preconditioner can be quite effective when the $\bm{G}$ has rapidly decaying eigenvalues and $r=\mathcal{O}(1)$, but the general case requires setting $r=\mathcal{O}(k)$, which is comparatively very expensive. A block-Lanczos-based preconditioner led to slow convergence in the numerical experiments of [38], suggesting that the approach is unlikely to be competitive with KRILL for challenging problems.
KRILL
In this paper, we recommend using a more accurate approximation,
@@
where $\bm{\Phi}\in{}^{d\times N}$ is a sparse random sign embedding (see Algorithm 5 or [15]). With this approximation, we can guarantee the effectiveness of the KRILL preconditioner $\bm{P}=\bm{\widehat{G}}+\mu\bm{A}(\mathsf{S},\mathsf{S})$, regardless of the number of centers $k$ and the regularization $\mu$.
KRILL is inspired by the sketch-and-precondition approach for solving overdetermined least-squares problems [29, §10.5]. In sketch-and-precondition, we use a random embedding to approximate the matrix appearing in the normal equations, and the approximate matrix serves as a preconditioner for solving the least-squares problem. Sketch-and-precondition was first proposed in [42] and later refined in [4, 33, 14, 26, 31]. It has been applied with several different embeddings [29, §§8–9], but the empirical comparisons of [19, Fig. 1] suggest the sparse sign embedding is the most efficient.
KRILL differs from classical presentations of sketch-and-precondition because the embedding is applied to the Gram matrix term $\bm{A}(:,\mathsf{S})\bm{A}(\mathsf{S},:)$ but not the regularization term $\mu\bm{A}(\mathsf{S},\mathsf{S})$. This idea of “partial sketching” appears in [39, 34], and it improves the accuracy of the approximation. Notwithstanding this difference, KRILL is motivated by the same ideas and its analysis follows the same patterns as classical sketch-and-precondition.
@@ 4 Case studies
In this section, we apply our preconditioning strategies to two scientific applications. In Section 4.1, we apply full-data KRR with RPCholesky preconditioning to predict the chemical properties of a wide range of molecules. In Section 4.2, we apply restricted KRR with KRILL preconditioning to distinguish exotic particle collisions from a background process.
@@ 4.1 HOMO energy prediction
A recent journal article [49], which was selected as an editor’s pick in the Journal of Chemical Physics, describes the difficulties in applying KRR to the QM9 data set. The authors aim to predict the highest-occupied-molecular-orbital (HOMO) energy. In their simplest approach, they make HOMO predictions based on the Coulomb matrix representation of molecules, which is a set of features based on the distances between the atomic nuclei and the nuclear charges. After defining the features, they standardize the features and apply an $\ell_{1}$ Laplace kernel
@@
Given the moderate size of the QM9 data set ($N=1.3\times 10^{5}$), it takes just a few seconds to load the data and form individual columns of the kernel matrix. However, on a laptop-scale computer (64GB RAM), there is not enough working memory to store or factorize the complete kernel matrix. The authors of [49] address these computational constraints by randomly subsampling $N=32,000$ or fewer data points and applying full-data KRR to the random sample. In their Fig. 8, they show that the predictive accuracy improves by a factor of 2×2$\times$2 × as they increase the size of the data from $N=1,000$ to $N=32,000$ data points. Even with $N=32,000$ data points, the predictive accuracy does not saturate, and thus computational constraints are limiting the the predictive accuracy of the KRR model.
@@
The results are presented in Figures 7, 8, and 9.
@@
Restricted KRR is a relatively cheap approach for predicting HOMO energies with the QM9 data set, since we can store the $k\times k$ preconditioner and $N\times k$ kernel submatrix in working memory with up to $k=50,000$ centers. Figure 7 shows that restricted KRR is more accurate than randomly subsampling the data and applying full-data KRR to the random sample. Still, the predictive accuracy increases with the number of centers $k$ and does not saturate even when $k=N/2$.
Full-data KRR is the most accurate approach for predicting HOMO energies, but it relies on repeated matrix–vector multiplication with the full kernel matrix. Since the kernel matrix is too large to store in 64GB of working memory, we can only store one block of the matrix at a time and perform the multiplications in a block-wise fashion. Each multiplication requires tens of minutes of computing, due to the cost of evaluating the kernel matrix each time the entries are needed. The training of the full-data KRR model is very slow unless we find a preconditioner that controls the total number of CG iterations.
@@
Figure 8 compares several preconditioning strategies for full-data KRR. The worst strategy is unpreconditioned CG, which leads to no perceivable convergence over the first $100$ iterations. Uniform and greedy Nyström preconditioning (with approximation rank $r=10^{3}$) improve the convergence speed, but they still require $100$ or more iterations for the predictive accuracy to saturate. RPCholesky (also with $r=10^{3}$) is the fastest method, leading to converged accuracy after $60$ iterations.
@@
Figure 9 shows that we can reduce the total number of CG iterations even further, by a factor of 5×5$\times$5 ×, if we increase the RPCholesky approximation rank from $r=10^{3}$ to $r=10^{4}$. However, we also need to consider the cost of preparing the preconditioner, which requires $\mathcal{O}(Nr^{2})$ operations. To achieve an appropriate balance, we recommend tuning the value of $r$ based on computational constraints, and $r$ should typically be small enough that the preconditioner can be stored in working memory. For example, an intermediate value of $10^{3}<r<10^{4}$ would be most appropriate for the HOMO energy prediction problem given our computing architecture.
Figure 9 shows a curious trend in which RPCholesky performs similarly to uniform pivot selection for a small approximation rank ($r<500$) and performs similarly to greedy pivot selection for a large approximation rank ($r>5000$). It is an open question to identify the problems and approximation ranks for which uniform or greedy pivot selection are as successful as RPCholesky.
@@ 4.2 Exotic particle detection
The Large Hadron Collider (LHC) is a particle accelerator that is being used to search for exotic particles not included in the standard model of physics. Each second, the LHC collides roughly one billion proton–proton pairs and produces one petabyte of data. Yet, exotic particles are believed to be produced in fewer than one per billion collisions, so only a small fraction of the data is relevant and the LHC is increasingly relying on machine learning to identify the most important observations for testing exotic particle models [40].
To help train machine learning algorithms, Baldi, Sadowsky, and Whiteson [7] have produced the SUSY data set ($N=5\times 10^{6}$) from Monte Carlo simulations of the decay processes of standard model bosons and supersymmetric particles not included in the standard model. Following [44, 45, 27], we will apply restricted KRR to the SUSY data to distinguish between the bosonic and supersymmetric decay processes.
For our KRR application, we split the data into $N=4.5\times 10^{6}$ training data points and $M=5\times 10^{5}$ testing data points, and we select $k=10^{4}$ data centers uniformly at random. Then we standardize the data features and apply a squared exponential kernel (2.3) with bandwidth $\sigma=4$. The parameters $k$ and $\sigma$ are the same ones used in previous work [44, 45], but we use a smaller regularization $\mu/N=2\times 10^{-10}$ (compared to $\mu/N=10^{-6}$ in [44, 45, 27]) which improves the test error from $19.6\%$ to $19.5\%$. We apply KRILL, FALKON, and unpreconditioned CG to solve the restricted KRR equations with these parameter choices. We terminate after forty iterations or when the relative residual falls below $\varepsilon=10^{-4}$, leading to the results in Figures 10 and 11.
@@
Figure 10 demonstrates the superior performance of KRILL, which reaches a test error of $19.5\%$ after just four iterations. Meanwhile, unpreconditioned CG fails to achieve such a low test error after forty iterations, and FALKON converges even more slowly than unpreconditioned CG. One reason for the slow convergence of FALKON is the small regularization parameter $\mu/N=2\times 10^{-10}$, which is a known failure mode for the preconditioner [27, Sec. 3]. Indeed, a previous application [45] of FALKON to the SUSY data set with a larger regularization $\mu/N=10^{-6}$ achieved test error convergence more quickly, in just twenty iterations (still five times slower than KRILL). A major advantage of KRILL is the reliable performance for the full range of $\mu$ values, which empowers users to choose the regularization that gives the lowest test error.
@@
Last, Figure 11 evaluates the total computation time needed to form the preconditioner and solve the restricted KRR equations up to a relative residual of $\varepsilon=10^{-4}$. For this figure, we randomly subsample $N=5\times 10^{5}$ data points so that the kernel submatrix fits in 64GB of working memory, and we compare KRILL, FALKON, and the direct method of forming and inverting $\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})$. We find that KRILL is the dominant algorithm once the number of centers reaches $k=1250$. KRILL outperforms the direct method because of the superior $\mathcal{O}((N+k^{2})k)$ computational cost, and KRILL outperforms FALKON because it converges in roughly 4×4$\times$4 × fewer CG iterations. As an advantage, the direct method is solving the restricted KRR equations up to machine precision, whereas KRILL and FALKON are only solving the equations up to a relative residual of $\varepsilon=10^{-4}$. Nonetheless, the test error plateaus at a threshold of $\varepsilon=10^{-4}$ in all our experiments.
@@ 5 Theoretical results
In this section, we prove our main theoretical results, Theorems 2.2 and 2.3.
5.1 Proof of Theorem 2.2
@@
$\kappa\bigl{(}\bm{P}^{-1/2}\bigl{(}\bm{A}+\mu\mathbf{I}\bigr{)}\bm{P}^{-1/2}\bigr{)}\leq 1+\operatorname{tr}\bigl{(}\bm{A}-\bm{\widehat{A}}\bigr{)}/\mu.$ (5.1)
To that end, observe that $\bm{P}\preceq\bm{A}+\mu\mathbf{I}$, because the Nyström approximation $\bm{\widehat{A}}$ is bounded from above by $\bm{A}$. By conjugation with $\bm{P}^{-1/2}$, we obtain
$\mathbf{I}\preceq\bm{P}^{-1/2}(\bm{A}+\mu\mathbf{I})\bm{P}^{-1/2}$
@@
$\bm{P}^{-1/2}(\bm{A}+\mu\mathbf{I})\bm{P}^{-1/2}=\mathbf{I}+\bm{P}^{-1/2}\big{(}\bm{A}-\bm{\widehat{A}}\big{)}\bm{P}^{-1/2}.$
Since the spectral norm is submultiplicative and equals the largest eigenvalue of a psd matrix, we deduce that
$\lambda_{\rm max}\bigl{(}\bm{P}^{-1/2}(\bm{A}+\mu\mathbf{I})\bm{P}^{-1/2}\bigr{)}$ $=1+\lambda_{\rm max}\bigl{(}\bm{P}^{-1/2}\big{(}\bm{A}-\bm{\widehat{A}}\big{)}\bm{P}^{-1/2}\bigr{)}$
$\leq 1+\left\|\smash{\bm{P}^{-1/2}}\right\|^{2}\left\|\smash{\bm{A}-\bm{\widehat{A}}}\right\|=1+\left\|\smash{\bm{P}^{-1}}\right\|\left\|\smash{\bm{A}-\bm{\widehat{A}}}\right\|.$
Last, using the fact that $\lVert\bm{A}-\bm{\widehat{A}}\rVert\leq\operatorname{tr}\big{(}\bm{A}-\bm{\widehat{A}}\big{)}$ and $\lambda_{\rm min}(\bm{P})\geq\mu$, we find
$\lambda_{\rm max}\bigl{(}\bm{P}^{-1/2}(\bm{A}+\mu\mathbf{I})\bm{P}^{-1/2}\bigr{)}\leq 1+\operatorname{tr}\big{(}\bm{A}-\bm{\widehat{A}}\big{)}/\mu.$ (5.3)
@@
As our next step, we apply the main RPCholesky error bound [12, Thm. 3.1] with rank $r=\text{rank}_{\mu}(\bm{A})$ and approximation accuracy $21+\varepsilon=$ 2, which guarantees
$\mathbb{E}\operatorname{tr}\bigl{(}\bm{A}-\bm{\widehat{A}}\bigr{)}\leq 2\sum\nolimits_{i>\text{rank}_{\mu}(\bm{A})}\lambda_{i}(\bm{A})$ (5.4)
@@
The right-hand side of (5.4) is bounded by $2\mu$ because of Definition 2.1 of the $\mu$-tail rank of $\bm{A}$. Hence, combining (5.4) with the condition number bound (5.1) guarantees
$\operatorname{\mathbb{E}}\kappa\Bigl{(}\bm{P}^{-1/2}\big{(}\bm{A}+\mu\mathbf{I}\big{)}\bm{P}^{-1/2}\Bigr{)}\leq 3.$
@@
$\frac{\lVert\bm{\beta}^{(t)}-\bm{\beta}\rVert_{\bm{A}+\mu\mathbf{I}}}{\lVert\bm{\beta}\rVert_{\bm{A}+\mu\mathbf{I}}}\leq 2\Biggl{(}\frac{\sqrt{3/\delta}-1}{\sqrt{3/\delta}+1}\Biggr{)}^{t}\leq 2\mathrm{e}^{-2t\sqrt{\delta/3}}\leq 2\mathrm{e}^{-t\sqrt{\delta}}$
@@
5.2 Proof of Theorem 2.3
Cohen [15, Thm. 4.2] showed that a sparse sign embedding with parameters $d=\mathcal{O}(k\log(k/\delta))$ and $\zeta=\mathcal{O}(\log(k/\delta))$ satisfies the following oblivious subspace embedding property: for any $k$-dimensional subspace $\mathsf{X}\subseteq{}^{N}$, there is probability at least $1-\delta$ that

### v5

@@ 3.2 Preconditioners for restricted KRR
where $\bm{\widehat{G}}\in\real^{k\times k}$ approximates the Gram matrix $\bm{G}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})$.
FALKON, the current state-of-the-art.
@@
These entries are the same in expectation:
@@
The main limitation of FALKON-type preconditioners is the Monte Carlo approximation error. To account for the Monte Carlo error, the available analyses [47, 48] assume a large number of centers $k=\Omega(\sqrt{N})$ and a large regularization $\mu=\Omega(\sqrt{N})$. When these assumptions are satisfied, FALKON-type preconditioners solve the restricted KRR equations in $\mathcal{O}((N+k^{2})k)$ operations. However, the assumptions on $k$ and $\mu$ may not be satisfied in practice. For example, $\mu$ is often small when chosen by cross-validation, and [47] contains examples of $\mu$ values up to five orders of magnitude smaller than $\mu=\sqrt{N}$.
@@
1:Embedding dimension $d$, input dimension NN, sparsity level $\zeta$.
2:Embedding matrix $\bm{\Phi}\in\real^{d\times N}$ in sparse (e.g., CSR) format.
[…]
@@
[…]
KRILL.
In this paper, we recommend the approximation
@@
where $\bm{\Phi}\in\real^{d\times N}$ is a sparse random sign embedding (see Algorithm˜5 or [16]). With this approximation, we can guarantee the effectiveness of the KRILL preconditioner $\bm{P}=\bm{\widehat{G}}+\mu\bm{A}(\mathsf{S},\mathsf{S})$, regardless of the number of centers $k$ and the regularization $\mu$.
KRILL is inspired by the sketch-and-precondition approach for solving overdetermined least-squares problems [33, §10.5]. In sketch-and-precondition, we use a random embedding to approximate the matrix appearing in the normal equations, and the approximation serves as a preconditioner for solving the least-squares problem. Sketch-and-precondition was first proposed in [45] and later refined in [4, 37, 15, 29, 35]. It has been applied with several different embeddings [33, §§8–9], but the empirical comparisons of [21, Fig. 1] suggest the sparse sign embedding is most efficient.
KRILL differs from classical presentations of sketch-and-precondition because the embedding is applied to the Gram matrix term $\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})$ but not the regularization term $\mu\bm{A}(\mathsf{S},\mathsf{S})$. This idea of “partial sketching” appears in [41, 38], and it improves the accuracy of the approximation. Notwithstanding this difference, KRILL is motivated by the same ideas as classical sketch-and-precondition and its analysis follows the same patterns.
@@ 4 Case studies
In this section, we apply our preconditioning strategies to two scientific problems. In Section˜4.1, we apply full-data KRR with RPCholesky preconditioning to predict the chemical properties of a wide range of molecules. In Section˜4.2, we apply restricted KRR with KRILL preconditioning to distinguish exotic particle collisions from a background process.
@@ 4.1 HOMO energy prediction
A recent journal article [53], which was selected as an editor’s pick in the Journal of Chemical Physics, describes the difficulties in applying KRR to the QM9 data set. The authors aim to predict the highest-occupied-molecular-orbital (HOMO) energy. In their simplest approach, they make HOMO predictions based on the Coulomb matrix representation of molecules, which is a set of features based on the distances between the atomic nuclei. After defining the features, they standardize the features and apply an $\ell_{1}$ Laplace kernel
@@
Given the moderate size of the QM9 data set ($N=1.3\times 10^{5}$), it takes just a few seconds to load the data and form individual columns of the kernel matrix. However, on a laptop-scale computer (64GB RAM), there is not enough working memory to store or factorize the complete kernel matrix. The authors of [53] address these computational constraints by randomly subsampling $N=32,000$ or fewer data points and applying full-data KRR to the random sample. In their Fig. 8, they show that the predictive accuracy improves by a factor of $2\times$ as they increase the size of the data from $N=1,000$ to $N=32,000$ data points. Even with $N=32,000$ data points, the predictive accuracy does not saturate, and thus computational constraints are limiting the the predictive accuracy of the KRR model.
@@
The results are presented in Figures˜7, 8, and 9.
@@
Restricted KRR is a relatively cheap approach for predicting HOMO energies with the QM9 data set, since we can store the $k\times k$ preconditioner and $N\times k$ kernel submatrix in working memory with up to $k=50,000$ centers. Figure˜7 shows that restricted KRR is more accurate than randomly subsampling the data and applying full-data KRR to the random sample. Still, the predictive accuracy increases with the number of centers $k$ and does not saturate even when $k=N/2$.
Full-data KRR is the most accurate approach for predicting HOMO energies, but it relies on repeated matrix–vector multiplications with the full kernel matrix. Since the kernel matrix is too large to store in 64GB of working memory, we can only store one block of the matrix at a time and perform the multiplications in a block-wise fashion. Each multiplication requires tens of minutes of computing, due to the cost of evaluating the kernel matrix each time the entries are needed. The training of the full-data KRR model is very slow unless we find a preconditioner that controls the total number of CG iterations.
@@
Figure˜8 compares several preconditioning strategies for full-data KRR. The worst strategy is unpreconditioned CG, which leads to no perceivable convergence over the first $100$ iterations. Uniform and greedy Nyström preconditioning (with an approximation rank $r=10^{3}$) improve the convergence speed, but they still require $100$ or more iterations for the predictive accuracy to saturate. RPCholesky (also with $r=10^{3}$) is the fastest method, leading to converged accuracy after $60$ iterations.
@@
Figure˜9 shows that we can reduce the total number of CG iterations even further, by a factor of $5\times$, if we increase the RPCholesky approximation rank from $r=10^{3}$ to $r=10^{4}$. However, we also need to consider the cost of preparing the preconditioner, which requires $\mathcal{O}(Nr^{2})$ operations. To achieve an appropriate balance, we recommend tuning the value of $r$ based on computational constraints, and $r$ should typically be small enough that the preconditioner can be stored in working memory. For example, an intermediate value of $10^{3}<r<10^{4}$ would be most appropriate for the HOMO energy prediction problem given our computing architecture.
Figure˜9 also reveals that RPCholesky performs similarly to uniform pivot selection for a small approximation rank ($r<500$) and performs similarly to greedy pivot selection for a large approximation rank ($r>5000$). It remains an open problem to identify the problems and approximation ranks for which uniform or greedy pivot selection are as successful as RPCholesky.
@@ 4.2 Exotic particle detection
The Large Hadron Collider (LHC) is a particle accelerator that is used to search for exotic particles not included in the standard model of physics. Each second, the LHC collides roughly one billion proton–proton pairs and produces one petabyte of data. Yet, exotic particles are believed to be produced in fewer than one per billion collisions, so only a small fraction of the data is relevant. The LHC uses machine learning to identify the most important observations for testing exotic particle models [42].
To help train machine learning algorithms, Baldi, Sadowsky, and Whiteson [7] produced the SUSY data set ($N=5\times 10^{6}$), derived from Monte Carlo simulations of the decay processes of standard model bosons and supersymmetric particles not included in the standard model. Following [47, 48, 30], we will apply restricted KRR to the SUSY data to distinguish between the bosonic and supersymmetric decay processes.
For our KRR application, we split the data into $N=4.5\times 10^{6}$ training data points and $M=5\times 10^{5}$ testing data points, and we select $k=10^{4}$ data centers uniformly at random. Then we standardize the data features and apply a squared exponential kernel (2.3) with bandwidth $\sigma=4$. The parameters $k$ and $\sigma$ are the same ones used in previous work [47, 48], but we use a smaller regularization $\mu/N=2\times 10^{-10}$ (compared to $\mu/N=10^{-6}$ in [47, 48, 30]) which improves the test error from $19.6\%$ to $19.5\%$. We apply KRILL, FALKON, and unpreconditioned CG to solve the restricted KRR equations with these parameter choices. We terminate after forty iterations or when the relative residual falls below $\varepsilon=10^{-4}$, leading to the results in Figures˜10 and 11.
@@
Figure˜10 demonstrates the superior performance of KRILL, which reaches a test error of $19.5\%$ after just four iterations. Meanwhile, unpreconditioned CG fails to achieve such a low test error after forty iterations, and FALKON converges even more slowly than unpreconditioned CG. One reason for the slow convergence of FALKON is the small regularization parameter $\mu/N=2\times 10^{-10}$ [30, Sec. 3]. Indeed, a previous application [48] of FALKON to the SUSY data set with a larger regularization $\mu/N=10^{-6}$ achieved test error convergence more quickly, in just twenty iterations (still five times slower than KRILL). A major advantage of KRILL is the reliable performance for the full range of $\mu$ values, which empowers users to choose the regularization that gives the lowest test error.
@@
Last, Figure˜11 evaluates the total computation time needed to form the preconditioner and solve the restricted KRR equations up to a relative residual of $\varepsilon=10^{-4}$. For this figure, we randomly subsample $N=5\times 10^{5}$ data points so the kernel submatrix fits in 64GB of working memory, and we compare KRILL, FALKON, and the direct method of forming and inverting $\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})$. We find that KRILL is the dominant algorithm once the number of centers reaches $k=1250$. KRILL outperforms the direct method because of the superior $\mathcal{O}(Nk\log k+k^{3})$ computational cost, and KRILL outperforms FALKON because it converges in roughly $4\times$ fewer CG iterations. As an advantage, the direct method solves the restricted KRR equations up to machine precision, whereas KRILL and FALKON only solve the equations up to a relative residual of $\varepsilon=10^{-4}$. Nonetheless, the test error plateaus at a threshold of $\varepsilon=10^{-4}$ in all our experiments.
@@ 5 Theoretical results
In this section, we prove our main theoretical results, Theorems˜2.2 and 2.3.
5.1 Proof of Theorem˜2.2
@@
[…]
$2\mathbf{I}\preceq\bm{P}^{-1/\penalty 502}(\bm{A}+\mu\mathbf{I})\bm{P}^{-1/\penalty 502}$
@@
[…]
Since the spectral norm is submultiplicative and it equals the largest eigenvalue of a psd matrix, we deduce that
[…]
@@
As our next step, we apply the main RPCholesky error bound [13, Thm. 3.1] with rank $r=\text{rank}_{\mu}(\bm{A})$ and approximation accuracy $1+\varepsilon=2$, which guarantees
$\mathbb{E}\operatorname{tr}\bigl(\bm{A}-\bm{\widehat{A}}\bigr)\leq 2\sum_{i>\text{rank}_{\mu}(\bm{A})}\lambda_{i}(\bm{A})$ (5.4)
@@
The right-hand side of (5.4) is bounded by $2\mu$ because of Definition˜2.1 of the $\mu$-tail rank of $\bm{A}$. Hence, combining (5.4) with the condition number bound (5.1) guarantees
[…]
@@
[…]
@@
5.2 Proof of Theorem˜2.3
Cohen [16, Thm. 4.2] showed that a sparse sign embedding with parameters $d=\mathcal{O}(k\log(k/\delta))$ and $\zeta=\mathcal{O}(\log(k/\delta))$ satisfies the following oblivious subspace embedding property: for any $k$-dimensional subspace $\mathsf{X}\subseteq\real^{N}$, there is probability at least $1-\delta$ that